OpenStudy (anonymous):

A circle has a diameter of 8 millimeters. What are the circumference and area of the circle? Use three point one four for pi. Round to the nearest tenth.

OpenStudy (anonymous):

Circumference: 25.1 mm Area: 50.2 mm To find the circumference, remember to use th formula: \[C = \pi x d \] (which can also be \[C = \pi x 2r\] ) To find the area, remember to use the formula: \[A = \pi x r^2\]

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

Each year, Terrell buys a plane ticket to visit his grandmother for a week. Last year, he paid one hundred thirty dollars. This year, he paid one hundred sixty-nine dollars. What was the percent increase in the cost of his plane ticket? Show your work.